A and B are two events and P (A) = 0.6, P (B) = 0.7 and P (A or B) = 0.9
Find P (not A and B)

Found 2 solutions by Boreal, greenestamps:
Answer by Boreal(15235)

P(A or B)=P(A)+P(B)-P(A and B)
0.9=0.6+0.7-0.4
Therefore since P(A and B) is 0.4, P not (A and B) is 0.6.

Answer by greenestamps(13198)

I read the problem differently than the other tutor.

From the given information, he correctly found that P(not(A AND B)) = 0.6.

As I read the problem, what we are supposed to find is P(not A AND B); that probability is P(B)-P(A AND B) = 0.7-0.4 = 0.3.